OpenStudy (anonymous):

Can someone help me? Solve the following system of equations x+5y=8 (1) x=5-5y (2)

OpenStudy (anonymous):

x+ 5y = 8 x + 5y = 5 So you have an inconsistent system of equations as x + 5y cannot be both values. Hope this helps. If it does, please choose good answer. Visit my website at http://www.tutorsean.net for free math training videos.

OpenStudy (anonymous):

yes he is right

OpenStudy (anonymous):

so as an ordered pair that would be (8,5)?

OpenStudy (anonymous):

graphically, the two lines will never intersect

OpenStudy (anonymous):

no, there in no ordered pair that satisfies both of these systems, theu theya re inconsistent

OpenStudy (anonymous):

so there is no solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

why is there no ordered pair that satisfys the solution?

OpenStudy (anonymous):

Because you have an inconsistent system. x cannot be 8 and 5 at the same time for example. If you are to look at this from a graphical perspective, let's take two answers that will satisfy the equations. Let x = 3 and y = 1. Then equation 1 is true, but equation 2 works out to be: 3 = 5 - 5(1) which gives us 3 = 0, which is not possible. Hope this helps. If it does, please choose good answer. Visit my website at http://www.tutorsean.net for free math training videos.